The following data shows the number of hours that 8 hospital patients slept following the administration of a certain anesthetic, the following 90% confidence interval for the average hours slept following the administration of the anesthetic was produced, (7.46, 9.88). What was the sample standard deviation?
Answer #1

E=9.88-8.67 = 1.21

t critical value =1.895...............by using t table or Excel =TINV(0.1,7)

Margin of error

Substitute here knwon values now

=> s=1.806

Sample standard deviation